### v1

# JanusDNA: A Powerful Bi-directional Hybrid DNA Foundation Model

Abstract
Large language models (LLMs) have revolutionized natural language processing and are increasingly applied to other sequential data types, including genetic sequences. However, adapting LLMs to genetics presents significant challenges. Capturing complex genomic interactions requires modeling long-range global dependencies within DNA sequences, where interactions often span over 10,000 base pairs, even within a single gene. This poses substantial computational demands under conventional model architectures and training paradigms. Additionally, traditional LLM training approaches are suboptimal for DNA sequences: autoregressive training, while efficient for training, only supports unidirectional sequence understanding. However, DNA is inherently bidirectional. For instance, bidirectional promoters regulate gene expression in both directions and govern approximately 11% of human gene expression. Masked language models (MLMs) enable bidirectional understanding. However, they are inefficient since only masked tokens contribute to loss calculations at each training step. To address these limitations, we introduce JanusDNA, the first bidirectional DNA foundation model built upon a novel pretraining paradigm, integrating the optimization efficiency of autoregressive modeling with the bidirectional comprehension capability of masked modeling. JanusDNA’s architecture leverages a Mamba-Attention Mixture-of-Experts (MoE) design, combining the global, high-resolution context awareness of attention mechanisms with the efficient sequential representation learning capabilities of Mamba. The MoE layers further enhance the model’s capacity through sparse parameter scaling, while maintaining manageable computational costs. Notably, JanusDNA can process up to 1 million base pairs at single-nucleotide resolution on a single 80GB GPU using its hybrid architecture. Extensive experiments and ablation studies demonstrate that JanusDNA achieves new state-of-the-art performance on three genomic representation benchmarks. Remarkably, JanusDNA surpasses models with 250x more activated parameters, underscoring its efficiency and effectiveness. Code available at https://github.com/Qihao-Duan/JanusDNA.
1 Introduction
Modeling the language of DNA is crucial for investigating its biological function, offering potential advancements in understanding genotype-phenotype associations, disease diagnosis, and new drug development benegas2025genomic . Large language models (LLMs) have demonstrated remarkable success in large-scale nonlinear representation learning for natural language processing (NLP) tasks, such as text generation, translation, and summarization liu2024deepseek ; bai2023qwen ; achiam2023gpt . This success has inspired researchers to explore LLM applications in other domains hu2025context ; jia2025seeing , including bioinformatics liu2025drbioright . However, applying general LLMs directly to DNA sequence data is challenging. DNA sequences are represented as series of nucleotides, lacking clear semantic meaning and posing challenges for interpretation by LLMs. Additionally, non-coding regions in DNA can be remotely related to coding regions both upstream and downstream, necessitating models capable of processing long-range dependencies while maintaining bidirectional understanding. These factors make it challenging to apply LLMs directly to DNA sequence data without significant modifications or adaptations caduceus . Some models have been developed specifically for DNA sequence representation dnabert ; dnabert2 ; nucleotidebenchmark . However, these models still face some limitations.
Current limitations
(1) Limited Sequence Length and Low Resolution: Capturing complex genomic interactions requires modeling long-range dependencies within DNA sequences, where interactions can span over 10,000 base pairs even within a single gene nemsick2024molecular . This necessitates a model that can effectively handle long-range dependencies and relationships within the sequence. However, many current models solely rely on global attention mechanisms inspired by their superior success in natural language applications jelassi2024repeat ; pantazopoulos2024shaking , yet they often struggle to effectively process long genomic sequences and uncover meaningful long-range interactions nucleotidebenchmark . K-mer tokenization is frequently used to expand the context window of DNA sequence models dnabert ; dnabert2 . However, this method introduces a trade-off between sequence length and resolution hyenadna , potentially leading to the loss of crucial information, especially in cases where single nucleotide polymorphisms (SNPs) are essential for understanding gene function. (2) Unidirectional Understanding: Many genomic processes are influenced by bidirectional interactions, with essential regulatory elements located both upstream and downstream of key genomic regions. For instance, gene expression is a complex process heavily influenced by regulatory sequences situated on either side of a gene. Accurately modeling these bidirectional interactions is crucial for understanding genomic function and making precise predictions, imposing significant demands on model architecture and training strategies caduceus . However, some existing SSM-based decoder-only models hyenadna ; evo ; evo2 are primarily unidirectional or limited in their capacity to effectively understand bidirectional context, thus constraining their ability to comprehensively capture these complex interactions. (3) Low Training Efficiency: Long-range modeling and bidirectional understanding of DNA sequences both require substantial computational resources and memory, especially for those requiring attention for a more global representation. Therefore, training efficiency significantly impacts the model’s performance, especially under limited computational resources. Most bidirectional models are trained using masked training paradigms, such as BERT devlin2018bert , which utilize only a small fraction of tokens (typically 15%) for loss calculation at each step. This limitation can hinder the model’s ability to learn effectively from the entire sequence within limited training steps, thereby requiring more training epochs to adequately cover the full training data.
Additionally, the masking process itself introduces extra computational overhead. In contrast, autoregressive (next-token prediction) training is more efficient, as nearly all tokens contribute to the loss at each training step, allowing the model to learn more effectively within a fixed number of steps as sequence length increases behnamghader2024llm2vec . However, it’s important to note that autoregressive models are inherently unidirectional, limiting their ability to incorporate bidirectional context.
In response to the aforementioned issues, we introduce JanusDNA [^1] Janus, the ancient Roman god of transitions and duality, is symbolized by two faces gazing in opposite directions., the first bidirectional DNA foundation model built upon a novel pretraining paradigm. Our architecture employs two principal strategies: (1) Hybrid Architecture: To achieve powerful global understanding while maintaining computational efficiency for long contexts, we integrate the strengths of state-space models (SSMs) mamba and Mixture-of-Experts (MoE) designs lieber2024jamba ; dai2024deepseekmoe into attention mechanisms pytorch_flexattention , enabling the model to effectively capture long-range global dependencies and complex interactions within DNA sequences. (2) Bidirectional Efficient Training: While preserving the bidirectional understanding of DNA sequences typically achieved through masked training, we significantly improve learning efficiency by computing the loss over all tokens in each training step like autoregressive modeling. Notably, JanusDNA is capable of processing up to 1 million base pairs at single-nucleotide resolution with global attention on a single 80GB GPU, making it suitable for large-scale understanding in genomic research. We evaluated JanusDNA on 35 diverse genomic tasks to showcase its superior global understanding as well as long-range representation ability.
In summary, our contributions are as follows:
  • We propose JanusDNA, a novel bidirectional DNA foundation model capable of capturing global long-range dependencies and interactions at single-nucleotide resolution.
  • We introduce an efficient Hybrid Mamba-Attention-MoE architecture designed for processing ultra-long genomic sequences within practical computational budgets.
  • We present Janus Modeling, a novel and efficient pretraining paradigm that effectively combines the strengths of autoregressive and masked modeling, facilitating effective global bidirectional learning.
  • We demonstrate state-of-the-art performance across diverse genomic benchmarks, outperforming significantly larger models. In particular, JanusDNA significantly surpasses the expert model Enformer on eQTL prediction tasks, despite having far fewer parameters.
2 Preliminary and Related Work
2.1 Large Language Model Pretraining Paradigms
Autoregressive Language Modeling (ALM)
is a generative pretraining paradigm in which the model predicts the next token in a sequence given all previous tokens. Trained on large corpora, the model learns statistical properties to generate coherent text. The training objective is:
$\mathcal{L}_{\text{CLM}}=-\sum_{t=1}^{T}\log P(x_{t}|x_{1},x_{2},\ldots,x_{t-1}),$ (1)
where $T$ is the sequence length, and $x_{t}$ denotes the token at position $t$. The model generates text by sampling from the learned probability distribution over the vocabulary at each time step hyenadna ; evo ; evo2 ; achiam2023gpt ; cui2024scgpt . Each token contributes to the overall loss, and the model minimizes the average loss across all tokens. However, to maximize generative performance cheng2024training , ALM is unidirectional, causing a limited ability to model bidirectional contexts, which is crucial for DNA sequence understanding benegas2025genomic ; consens2023transformers .
Masked Language Modeling (MLM)
is a non-causal pretraining paradigm where the model predicts masked tokens in a sequence using surrounding context. The training objective is:
$\mathcal{L}_{\text{MLM}}=-\sum_{i=1}^{N}\log P(x_{i}|x_{j_{1}},x_{j_{2}},\ldots,x_{j_{k}}),$ (2)
where $N$ is the total number of tokens, $x_{i}$ is the masked token, and $x_{j_{1}},x_{j_{2}},\ldots,x_{j_{k}}$ are the unmasked tokens. This approach enables the model to learn bidirectional representations, capturing dependencies in both directions benegas2025genomic ; cheng2024training ; consens2023transformers ; dnabert ; dnabert2 ; theodoris2023transfer ; nucleotidebenchmark ; benegas2023dna . However, MLMs mostly follow the BERT-style training paradigm, which masks a fixed percentage of tokens in the input sequence, e.g., 15% caduceus ; dnabert ; dnabert2 . This can lead to inefficiencies, as only a small fraction of the data is used for loss computation during each iteration. In contrast, autoregressive training paradigms take advantage of nearly the entire data, significantly improving training efficiency and overall performance.
A detailed review of related work on DNA language models is provided in Appendix A.
3 JanusDNA
Figure 1: The JanusDNA Architecture for Bidirectional DNA Modeling. JanusDNA employs a hierarchical bidirectional strategy to comprehensively model DNA sequences. (A) DNA, with its inherent double-stranded nature. (B) The model processes both the forward and reverse complement strands independently in parallel to capture complete biological context, with their embeddings subsequently combined for downstream tasks. (C) The core JanusDNA model architecture processes a single input strand using parallel left-to-right and right-to-left pathways. Each pathway consists of Mamba-FFN and Mamba-MoE layers for effective and efficient sequential encoding. (D) The MoE architecture enhances model capacity and specialization by dynamically and sparsely routing inputs to a subset of expert networks, enabling efficient computation and improved representation learning. (E) The Bidirectional Global Fusion mechanism, utilizing a specific attention mask, integrates the forward and backward representations from (C) to ensure that each nucleotide’s embedding is informed by its complete sequence context.
We propose Janus modeling, an efficient bidirectional training method with global attention, and JanusDNA, a powerful hybrid DNA foundation model.
As illustrated in Figure 1, JanusDNA processes bidirectional DNA sequences from both left-to-right and right-to-left using two independent stacks of Mamba and Mixture-of-Experts (MoE) layers. These stacks generate forward and backward representations independently, ensuring no information leakage. The two representations are then fused to create a unified representation that encapsulates bidirectional information. Each token position is predicted based on all upstream and downstream tokens. The following sections detail the efficient bidirectional training method with global attention – Janus Modeling, and the hybrid Mixture-of-Experts (MoE) architecture – JanusDNA.
3.1 Bidirectional Efficient Training
As discussed in Section 2, conventional pretraining paradigms face a trade-off: Masked Language Models (MLMs) offer bidirectional understanding but suffer from low training efficiency due to sparse loss signals, especially for those requiring global attention, while Autoregressive Models are efficient in training but inherently unidirectional. To overcome this, we introduce Janus modeling, a novel pretraining objective designed to achieve efficient, fully bidirectional sequence understanding with global attention, as illustrated in Figure 2.
Figure 2: Modeling Interpretation. Janus modeling treats each token as a target for loss calculation, enabling higher training efficiency compared to masked modeling by full sequence learning while keeping bidirectional context understanding.
The core idea of Janus modeling is to predict every token $x_{t}$ in a sequence of length $T$ using its complete bidirectional context, i.e., all tokens preceding $x_{t}$ ($x_{1},\ldots,x_{t-1}$) and all tokens succeeding $x_{t}$ ($x_{t+1},\ldots,x_{T}$). The training objective is therefore:
$\mathcal{L}_{\text{bidirectional}}=-\sum_{t=1}^{T}\log P(x_{t}|x_{1},\ldots,x_{t-1},x_{t+1},\ldots,x_{T})$ (3)
This objective ensures that every token contributes to the loss, maximizing training efficiency, while simultaneously demanding bidirectional comprehension.
To realize this objective, Janus modeling employs a two-stage process: independent context encoding followed by a global fusion mechanism:
Independent Context Encoding:
The input sequence $X=(x_{1},\ldots,x_{T})$ is processed by two parallel and independent stacks of layers (detailed in Section 3.2):
  • A forward pass processes the sequence from left-to-right, generating a sequence of hidden states $\mathcal{R}_{\text{fwd}}=(H^{F}_{1},H^{F}_{2},\ldots,H^{F}_{T})$. Each hidden state $H^{F}_{t}$ is a function of the input tokens $(x_{1},\ldots,x_{t})$ and primarily captures information about the left context of $x_{t}$.
    $H^{F}_{t}=\text{ForwardEncoder}(x_{1},\ldots,x_{t})$ (4)
  • A backward pass processes the sequence from right-to-left, generating a sequence of hidden states $\mathcal{R}_{\text{bwd}}=(H^{B}_{1},H^{B}_{2},\ldots,H^{B}_{T})$. Each hidden state $H^{B}_{t}$ is a function of the input tokens $(x_{T},\ldots,x_{t})$ and primarily captures information about the right context of $x_{t}$.
    $H^{B}_{t}=\text{BackwardEncoder}(x_{T},\ldots,x_{t})$ (5)
These two sets of representations, $\mathcal{R}_{\text{fwd}}$ and $\mathcal{R}_{\text{bwd}}$, are generated independently, ensuring no premature information leakage between past and future contexts before the explicit fusion step. The entire model is trained end-to-end using $\mathcal{L}_{\text{bidirectional}}$ from Equation 3.
Bidirectional Global Fusion:
To compute $P(x_{t}|x_{1},\ldots,x_{t-1},x_{t+1},\ldots,x_{T})$ for each $x_{t}$ as per Equation 3, the left-context information captured in $\mathcal{R}_{\text{fwd}}$ and the right-context information captured in $\mathcal{R}_{\text{bwd}}$ must be integrated. This is organically achieved via a global attention mechanism, specifically implemented with FlexAttention pytorch_flexattention for efficiency. The representations from both passes, $\mathcal{R}_{\text{fwd}}=(H^{F}_{1},\ldots,H^{F}_{T})$ and $\mathcal{R}_{\text{bwd}}=(H^{B}_{1},\ldots,H^{B}_{T})$, are concatenated to form a combined input sequence for the attention layer: $R_{\text{input}}=[H^{F}_{1},\ldots,H^{F}_{T},H^{B}_{1},\ldots,H^{B}_{T}]$. This $R_{\text{input}}$ sequence has a length of $2T$. The core of the fusion lies in a carefully designed attention mask, $\mathcal{M}_{ij}$, which dictates how tokens in $R_{\text{input}}$ can attend to each other. This mask ensures that the prediction for $x_{t}$ is based only on $H^{F}_{k}$ for $k<t$ and $H^{B}_{j}$ for $j>t$, preventing information leakage. The mask, also illustrated in Figure 1(E), is defined as:
$\mathcal{M}_{ij}=\begin{cases}q_{\text{idx}}\geq kv_{\text{idx}},&\text{if }kv_{\text{idx}}<T\text{ and }q_{\text{idx}}<T,\\ q_{\text{idx}}\leq kv_{\text{idx}},&\text{if }kv_{\text{idx}}\geq T\text{ and }q_{\text{idx}}\geq T,\\ kv_{\text{idx}}\geq T+q_{\text{idx}}+2,&\text{if }kv_{\text{idx}}\geq T\text{ and }q_{\text{idx}}<T,\\ q_{\text{idx}}\geq kv_{\text{idx}}+T+2,&\text{if }kv_{\text{idx}}<T\text{ and }q_{\text{idx}}\geq T.\end{cases}$ (6)
Here, $q_{\text{idx}}$ and $kv_{\text{idx}}$ are the 0-indexed indices of the query and key-value pairs within the $2T$-length $R_{\text{input}}$, respectively. $T$ is the original sequence length. The mask $\mathcal{M}_{ij}$ is a binary matrix where allowed attentions are 1 and disallowed are 0 (or $\infty$- ∞ after softmax). The first two cases handle causal attention within the $\mathcal{R}_{\text{fwd}}$ and $\mathcal{R}_{\text{bwd}}$ segments, respectively. The third and fourth cases manage the cross-attention between $\mathcal{R}_{\text{fwd}}$ and $\mathcal{R}_{\text{bwd}}$ segments, precisely controlling information flow to maintain the integrity of bidirectional prediction without future peeking relative to the token being predicted. The output of this attention mechanism provides a fused representation $H^{\text{final}}_{t}$ for each token $x_{t}$, which is then used to make the final prediction, optimizing $\mathcal{L}_{\text{bidirectional}}$.
This Janus modeling approach, as conceptually depicted in Figure 2 (C), enables each token to be a learning target informed by its full bidirectional context, thereby enhancing training efficiency compared to traditional MLMs (Figure 2 (B)) and overcoming the limitations of unidirectional models (Figure 2 (A)).
Empirical Validation of Training Efficiency
To empirically assess the learning efficiency of Janus modeling against conventional bidirectional approaches – masked modeling, we conducted a comparative experiment focused on last-token prediction. This task was chosen as it allows a direct comparison: both a standard masked language model and our Janus modeling approach can predict the final token $x_{T}$ given the preceding context $x_{1},\ldots,x_{T-1}$, ensuring a fair basis for evaluation.
For this demonstration, we configured two model variants: Janus Model: A single-layer bidirectional Mamba architecture equipped with a FlexAttention layer pytorch_flexattention for bidirectional fusion, utilizing the Janus-specific attention mask $\mathcal{M}_{ij}$ (Equation 6). This model predicts $x_{t}$ based on the fused bidirectional context of $x_{1},\ldots,x_{t-1},x_{t+1},\ldots,x_{T}$ as per the Janus modeling objective. Masked Language Model: As a baseline, we construct a comparable single-layer Mamba architecture followed by the same FlexAttention layer without an attention mask, trained using the conventional masked language modeling (MLM) objective, where a fraction of tokens (typically 15%) contribute to the loss.
Figure 3: Superior Learning Efficiency of Janus Modeling. Comparison of last-token prediction accuracy between Janus modeling and conventional masked modeling over 10k training steps. Janus modeling consistently achieves higher accuracy for the same model architecture and training duration, demonstrating its enhanced efficiency in learning from sequence data.
Both models were trained on the human reference genome (HG38) genome2009genome for 10,000 steps. Each training sample had a context length of 131,072 tokens, processed with a batch size of 1. We evaluated performance across three hidden dimensions: 32, 64, and 128, keeping other hyperparameters consistent. Pre-training for Janus models takes around 27 minutes per 1,000 steps, nearly twice as fast as masked models.
For evaluation, the masked model is given the full input sequence of $T$ tokens with the last token $x_{T}$ masked, while the Janus model only sees the first $x_{1},\ldots,x_{T-1}$ tokens. This evaluation is conducted on a test set containing 1,920 sequences.
The results, presented in Figure 3, clearly demonstrate that models trained with the Janus modeling method significantly outperformed those trained with the masked modeling approach in prediction accuracy, given the same number of training steps. This finding substantiates that Janus modeling is more effective at leveraging the DNA sequences for learning, thereby achieving superior training efficiency while maintaining robust bidirectional understanding.
3.2 Hybrid Mixture-of-Experts (MoE) Model
After introducing the bidirectional efficient training method, we developed a bidirectional backbone model to enhance sequence representation. Leveraging the efficient bidirectional fusion method, we propose JanusDNA, a hybrid model that integrates the strengths of Mamba, Attention, and MoE.
Architecture
As illustrated in Fig. 1, JanusDNA incorporates three primary components for sequence representation: Mamba, MoE, and Attention. Mamba, a state space model (SSM) mamba , efficiently encodes input sequences using high-dimensional parameters. Compared to traditional Transformer architectures, Mamba is more memory- and computationally efficient, making it particularly suitable for processing long DNA sequences.
The Mixture-of-Experts (MoE) architecture provides a scalable approach to significantly increase model capacity without proportionally increasing computational costs during training and inference Cai_2025 . In JanusDNA, MoE layers replace the feedforward network (FFN) layers in the Mamba blocks at a specific ratio following lieber2024jamba , achieving a balance between performance and efficiency. To ensure balanced utilization of the experts, an auxiliary loss is introduced switch_transformer , encouraging the model to distribute input data evenly across all experts. This auxiliary loss is computed based on the router logits, which represent the probabilities of selecting each expert for a given input, as shown in Eq. 7. Given $N$ experts indexed by $i=1$ to $N$ and a batch $\mathcal{B}$ with $T$ tokens, the auxiliary loss is computed as the scaled dot-product between vectors $f$ and $P$:
$\mathcal{L}_{\text{total}}=\alpha\cdot N\cdot\sum_{i=1}^{N}f_{i}\cdot P_{i}$ (7)
where $f_{i}$ is the fraction of tokens dispatched to expert $i$,
$f_{i}=\frac{1}{T}\sum_{x\in\mathcal{B}}1\{\text{argmax}\>p(x)=i\},$ (8)
and $P_{i}$ is the fraction of the router probability allocated to expert $i$,
$P_{i}=\frac{1}{T}\sum_{x\in\mathcal{B}}p_{i}(x).$ (9)
Here, $p_{i}(x)$ represents the probability of routing token $x$ to expert $i$, while $T$ is the total number of tokens in the batch $\mathcal{B}$. This auxiliary loss encourages balanced utilization of all experts, improving overall model performance.
Bidirectional sequences are processed through independent stacks of Mamba and MoE layers, each designed to enhance the model’s representational capacity. The Mamba layers efficiently capture local contextual dependencies within the sequence, leveraging their memory-efficient state-space modeling, while the MoE layers provide sparse scaling to enhance the model’s representational capacity without incurring proportional computational overhead.
The forward and backward representations generated by these layers are fused using FlexAttention pytorch_flexattention , an optimized attention mechanism that supports sparse masks with reduced memory consumption. This fusion enables the model to integrate both local and long-range global information streams, resulting in a comprehensive bidirectional representation for improved performance.
Reverse Complement
In the double-helix DNA structure, each strand contains semantically equivalent information, with the reverse complement (RC) of a strand oriented in the opposite direction and its bases complemented relative to the forward strand (A paired with T, and C paired with G) caduceus . However, recognizing both the forward and RC versions of non-palindromic motifs, such as GATA and TATC, poses a significant challenge, as it is akin to learning two distinct motifs zhou2022towards . To address this, we adopt a post-hoc reverse complement representation strategy alipanahi2015predicting . Specifically, the DNA sequence and its reverse complement are processed in parallel using the identical model. The resulting representation vectors are then pooled to form a unified, enriched representation as shown in Figure 1 (B). This approach enables the model to effectively learn from both the original and reverse complement sequences, improving its ability to capture intricate patterns and relationships within DNA sequences, thereby enhancing performance across various tasks.
4 Experiments
4.1 Pre-training on Human Reference Genome
To ensure a fair comparison with prior work, we pre-train our model on only the human reference genome (HG38 genome2009genome ) following the training setup described in caduceus . Specifically, we adopt single nucleotide-level tokenization to capture high-resolution input sequences and avoid overlooking critical DNA information that may be lost when using k-mer tokenization, commonly used in attention-based models hyenadna . Additionally, single nucleotide-level tokenization is employed to facilitate downstream research on Single Nucleotide Polymorphisms (SNPs). Please note that performance on downstream tasks depends on the model architecture, as well as the composition and diversity of the pretraining data nucleotidebenchmark ; evo ; evo2 ; dnabert2 . Here, we specifically focus only on the architecture, and thus use only the human reference genome for pretraining to ensure a fair comparison.
4.2 Downstream Tasks
We evaluate our model on three different benchmarks: Genomic Benchmark genomicbenchmark , Nucleotide Transformer Benchmark nucleotidebenchmark , and DNALONGBENCH chengdna . We follow all benchmark settings of Genomic Benchmark and Nucleotide Transformer Benchmark as described in caduceus . Accordingly, we adopt the reported results from caduceus as our reference. Considering the practical computational cost of sparse MoEs, we adjust the model’s hidden size to match or slightly reduce the number of activated parameters compared to the baseline caduceus , ensuring a fair comparison. As we introduce the model with a middle attention layer in the ablation experiments (Section C), we present results for models both with and without mid-attention on the Genomic Benchmark and Nucleotide Transformer Benchmark.
4.2.1 Genomic Benchmark
Table 1: Genomic Benchmarks. Top-1 accuracy ($\uparrow$) across 5-fold cross-validation (CV) for a supervised CNN baseline, pretrained HyenaDNA, Mamba NTP, Caduceus models, and $\sf JanusDNA$ models. Best values per task are bolded, second best are underlined. Error bars indicate the difference between the maximum and minimum values across 5 random seeds used for CV.
Models
activated param
CNN
HyenaDNA
Mamba
Caduceus
Ph
Caduceus
PS
JanusDNA
w/ mid-Attn
JanusDNA
w/o mid-Attn
Mouse Enhancers $0.715$±$\pm 0.087$ $0.780$±$\pm 0.025$ $0.743$±$\pm 0.054$ $0.754$±$\pm 0.074$ $0.793$±$\pm 0.058$ $0.752$±$\pm 0.041$ $0.750$±$\pm 0.031$
Coding vs. Intergenomic $0.892$±$\pm 0.008$ $0.904$±$\pm 0.005$ $0.904$±$\pm 0.004$ $0.915$±$\pm 0.003$ $0.910$±$\pm 0.003$ $0.913$±$\pm 0.001$ $0.914$±$\pm 0.002$
Human vs. Worm $0.942$±$\pm 0.002$ $0.964$±$\pm 0.002$ $0.967$±$\pm 0.002$ $0.973$±$\pm 0.001$ $0.968$±$\pm 0.002$ $0.972$±$\pm 0.001$ $0.971$±$\pm 0.001$
Human Enhancers Cohn $0.702$±$\pm 0.021$ $0.729$±$\pm 0.014$ $0.732$±$\pm 0.029$ $0.747$±$\pm 0.004$ $0.745$±$\pm 0.007$ $0.745$±$\pm 0.004$ $0.740$±$\pm 0.002$
Human Enhancer Ensembl $0.744$±$\pm 0.122$ $0.849$±$\pm 0.006$ $0.862$±$\pm 0.008$ $0.893$±$\pm 0.008$ $0.900$±$\pm 0.006$ $0.899$±$\pm 0.003$ $0.900$±$\pm 0.005$
Human Regulatory $0.872$±$\pm 0.005$ $0.869$±$\pm 0.012$ $0.814$±$\pm 0.211$ $0.872$±$\pm 0.011$ $0.873$±$\pm 0.007$ $0.877$±$\pm 0.002$ $0.870$±$\pm 0.006$
Human OCR Ensembl $0.698$±$\pm 0.013$ $0.783$±$\pm 0.007$ $0.815$±$\pm 0.002$ $0.828$±$\pm 0.006$ $0.818$±$\pm 0.006$ $0.825$±$\pm 0.002$ $0.825$±$\pm 0.003$
Human NonTATA Promoters $0.861$±$\pm 0.009$ $0.944$±$\pm 0.002$ $0.933$±$\pm 0.007$ $0.946$±$\pm 0.007$ $0.945$±$\pm 0.010$ $0.949$±$\pm 0.006$ $0.945$±$\pm 0.006$
We start with the Genomic Benchmark, which is a collection of 8 regulatory element classification tasks with sequence lengths mostly ranging from 200 to 500, and one up to 4,776. We take the hidden state embedding of the final layer and apply a pooling layer on sequences to obtain a fixed-length representation. We then apply a linear layer to map the representation to the number of classes for each task. We perform 5-fold cross-validation for each task using the same seed as caduceus . As shown in Table 1, our model achieves state-of-the-art performance on 3 out of 8 tasks, outperforming the previous best model, while the remaining tasks are close to the best model. Please note that this benchmark is already quite saturated, as shown in Table 1, and we do not expect improvements in pretraining to meaningfully improve benchmark performance further.
Table 2: Nucleotide Transformer Tasks. Performance ($\uparrow$) across 10-fold CV for Enformer, DNABERT-2, Nucleotide Transformer v2, HyenaDNA, Caduceus-PH, and $\sf JanusDNA$. Metrics vary by task: MCC for histone markers and enhancer annotation, F1-score for promoter annotation and splice site acceptor/donor, and accuracy for splice site “all”. Best values per task are bolded, second best are italicized. Given the disparity in model size, we also underline the best value among models with fewer than 2M activated parameters. Error bars indicate the difference between the maximum and minimum values across 10 random seeds used for CV.
$>$ 100M activated Param. Models $<$ 2M activated Param. Models
Enformer
DNABERT-2
NT-v2
HyenaDNA
Caduceus-Ph
janusDNA
w/ mid-attn
janusDNA
w/o-midattn
Histone Markers
H3 $0.719$±$\pm 0.048$ $0.785$±$\pm 0.033$ $0.784$±$\pm 0.047$ $0.779$±$\pm 0.037$ $0.815$±$\pm 0.048$ $0.821$±$\pm 0.021$ $0.824$±$\pm 0.012$
H3k14ac $0.288$±$\pm 0.077$ $0.516$±$\pm 0.028$ $0.551$±$\pm 0.021$ $0.612$±$\pm 0.065$ $0.631$±$\pm 0.026$ $0.665$±$\pm 0.034$ $0.685$±$\pm 0.016$
H3k36me3 $0.344$±$\pm 0.055$ $0.591$±$\pm 0.020$ $0.625$±$\pm 0.013$ $0.613$±$\pm 0.041$ $0.601$±$\pm 0.129$ $0.658$±$\pm 0.024$ $0.670$±$\pm 0.012$
H3k4me1 $0.291$±$\pm 0.061$ $0.511$±$\pm 0.028$ $0.550$±$\pm 0.021$ $0.512$±$\pm 0.024$ $0.523$±$\pm 0.039$ $0.563$±$\pm 0.041$ $0.571$±$\pm 0.018$
H3k4me2 $0.211$±$\pm 0.069$ $0.336$±$\pm 0.040$ $0.319$±$\pm 0.045$ $0.455$±$\pm 0.095$ $0.487$±$\pm 0.170$ $0.509$±$\pm 0.056$ $0.548$±$\pm 0.022$
H3k4me3 $0.158$±$\pm 0.072$ $0.352$±$\pm 0.077$ $0.410$±$\pm 0.033$ $0.549$±$\pm 0.056$ $0.544$±$\pm 0.045$ $0.605$±$\pm 0.030$ $0.629$±$\pm 0.022$
H3k79me3 $0.496$±$\pm 0.042$ $0.613$±$\pm 0.030$ $0.626$±$\pm 0.026$ $0.672$±$\pm 0.048$ $0.697$±$\pm 0.077$ $0.716$±$\pm 0.017$ $0.727$±$\pm 0.023$
H3K9ac $0.420$±$\pm 0.063$ $0.542$±$\pm 0.029$ $0.562$±$\pm 0.040$ $0.581$±$\pm 0.061$ $0.622$±$\pm 0.030$ $0.641$±$\pm 0.024$ $0.639$±$\pm 0.019$
H4 $0.732$±$\pm 0.076$ $0.796$±$\pm 0.027$ $0.799$±$\pm 0.025$ $0.763$±$\pm 0.044$ $0.811$±$\pm 0.022$ $0.809$±$\pm 0.021$ $0.816$±$\pm 0.008$
H4ac $0.273$±$\pm 0.063$ $0.463$±$\pm 0.041$ $0.495$±$\pm 0.032$ $0.564$±$\pm 0.038$ $0.621$±$\pm 0.054$ $0.637$±$\pm 0.060$ $0.653$±$\pm 0.034$
Regulatory Annotation
Enhancer $0.451$±$\pm 0.108$ $0.516$±$\pm 0.098$ $0.548$±$\pm 0.144$ $0.517$±$\pm 0.117$ $0.546$±$\pm 0.073$ $0.564$±$\pm 0.022$ $0.535$±$\pm 0.036$
Enhancer types $0.309$±$\pm 0.134$ $0.423$±$\pm 0.051$ $0.424$±$\pm 0.132$ $0.386$±$\pm 0.185$ $0.439$±$\pm 0.054$ $0.462$±$\pm 0.049$ $0.470$±$\pm 0.025$
Promoter: All $0.954$±$\pm 0.006$ $0.971$±$\pm 0.006$ $0.976$±$\pm 0.006$ $0.960$±$\pm 0.005$ $0.970$±$\pm 0.004$ $0.969$±$\pm 0.002$ $0.971$±$\pm 0.002$
     NonTATA $0.955$±$\pm 0.010$ $0.972$±$\pm 0.005$ $0.976$±$\pm 0.005$ $0.959$±$\pm 0.008$ $0.969$±$\pm 0.011$ $0.971$±$\pm 0.003$ $0.971$±$\pm 0.002$
     TATA $0.960$±$\pm 0.023$ $0.955$±$\pm 0.021$ $0.966$±$\pm 0.013$ $0.944$±$\pm 0.040$ $0.953$±$\pm 0.016$ $0.956$±$\pm 0.010$ $0.958$±$\pm 0.008$
Splice Site Annotation
All $0.848$±$\pm 0.019$ $0.939$±$\pm 0.009$ $0.983$±$\pm 0.008$ $0.956$±$\pm 0.011$ $0.940$±$\pm 0.027$ $0.963$±$\pm 0.022$ $0.960$±$\pm 0.009$
Acceptor $0.914$±$\pm 0.028$ $0.975$±$\pm 0.006$ $0.981$±$\pm 0.011$ $0.958$±$\pm 0.010$ $0.937$±$\pm 0.033$ $0.949$±$\pm 0.020$ $0.939$±$\pm 0.022$
Donor $0.906$±$\pm 0.027$ $0.963$±$\pm 0.006$ $0.985$±$\pm 0.022$ $0.949$±$\pm 0.024$ $0.948$±$\pm 0.025$ $0.947$±$\pm 0.015$ $0.936$±$\pm 0.014$
4.2.2 Nucleotide Transformer Tasks
Next, we evaluate our model on the Nucleotide Transformer tasks, which include 18 datasets covering histone marker prediction, regulatory annotation prediction, and splice site annotation prediction. Following the evaluation metrics outlined in nucleotidebenchmark , we perform 10-fold cross-validation for each task, adhering to the same experimental settings as caduceus .
As shown in Table 2, our model achieves state-of-the-art performance on 12 out of 18 tasks, surpassing previous models, including those with 250 times more activated parameters. While the promoter and splice site annotation tasks exhibit slightly weaker performance compared to the best larger model, this underscores the potential importance of training data scale and diversity for these specific tasks. For clarity, we present only the results of Caduceus-PH in the table due to space constraints, as Caduceus-PS performs slightly worse than Caduceus-PH.
4.2.3 DNA Long Range Benchmark
To further assess our model’s ability to capture long-range dependencies in DNA sequences, we evaluate it on the expression Quantitative Trait Loci (eQTL) prediction task from DNALONGBENCH chengdna . The eQTL task measures whether a nucleotide variant can influence the expression of a target gene based on the sequences of the gene and its surrounding regions.
Table 3: DNALongBench eQTL Tasks. The AUROC for expert model - Enformer, Caduceus-PH, and $\sf JanusDNA$. The best results are bolded.
Models
activated param
expert model
Caduceus-Ph
JanusDNA
w/o mid-Attn
Artery Tibial $0.741$ $0.690$ $0.803$
Adipose Subcutaneous $0.736$ $0.759$ $0.741$
Cells Cultured fibroblasts $0.639$ $0.690$ $0.771$
Muscle Skeletal $0.621$ $0.789$ $0.803$
Nerve Tibial $0.683$ $0.842$ $0.877$
Skin Not Sun Exposed Suprapubic $0.710$ $0.812$ $0.875$
Skin Sun Exposed Lower leg $0.700$ $0.692$ $0.706$
Thyroid $0.612$ $0.703$ $0.752$
Whole Blood $0.689$ $0.769$ $0.794$
Due to limited computational resources, we compare our model against the state-of-the-art DNA language model, Caduceus-PH caduceus , which is also trained with the same data scale for fair comparison, and the expert model for this task, Enformer enformer . The results for Enformer are taken directly from DNALONGBENCH chengdna . For Caduceus-PH, we fine-tune the official HuggingFace-released weights, which are pretrained on sequences of length 131k. Our model is pretrained and fine-tuned using the same setup and sequence length as Caduceus-PH to ensure a fair comparison.
As shown in Table 3, JanusDNA achieves the best performance on 8 out of 9 datasets, significantly outperforming the expert model and Caduceus-PH despite using fewer computational parameters.
5 Conclusion
Summary
In this work, we introduced a novel global modeling paradigm for bidirectional DNA sequence representation, combining the bidirectional capability of masked language modeling with the speed and optimization benefits of autoregressive approaches. We proposed JanusDNA, a Mamba MoE-based DNA foundation model with global attention that enhances genomic sequence understanding while maintaining low memory complexity, supporting the processing of up to 1 million base pairs (1 Mbp) on a single 80GB GPU. Experimental results demonstrate that JanusDNA outperforms HyenaDNA, Caduceus, and other Transformer-based models across a range of benchmark tasks and the expert model on eQTL tasks. By leveraging global attention mechanisms and efficient long-range sequence processing, JanusDNA offers a powerful framework for advancing research on long-range genomic interactions.
Limitations and Future Work
Although JanusDNA demonstrates high learning efficiency on a fixed data scale, current training is restricted to the human reference genome for fair architectural comparison. Expanding the corpus to include human genomic variants (e.g., 1000 Genomes Project) and non-human species (e.g., primates) could further boost modeling capacity and biological insight. JanusDNA also lacks integration of multimodal data such as epigenetic states (e.g., chromatin accessibility, histone modifications) and single-cell transcriptomic profiles, which are vital for resolving cell-type-specific regulation and predicting chromatin-influenced phenotypes. Future work will incorporate these modalities and explore functional roles of key genomic features (e.g., CTCF-mediated chromatin loops, enhancer RNAs, non-coding risk variants). Experimental validation (e.g., CRISPR, organoids) will prioritize therapeutic targets, while clinical collaborations will evaluate JanusDNA’s utility in personalized diagnostics and drug discovery. These efforts aim to evolve JanusDNA into a unified framework linking genome structure, epigenetic regulation, and disease mechanisms.
6 Acknowledgments
The authors acknowledge the Scientific Computing of the IT Division at the Charité - Universitätsmedizin Berlin and the Science Computing at Shenzhen Technology University for providing computational resources that have contributed to the research results reported in this paper.
Appendix A Related Work
A.1 DNA Language Models
Attention-based Models
Attention-based DNA language models, such as DNABERT [9], DNABERT2 [10], and Nucleotide Transformer [11], have demonstrated significant success in modeling DNA sequences. These models employ k-mer encoders to group consecutive nucleotide base pairs into single tokens, enabling efficient sequence representation. However, their global attention mechanisms restrict scalability to sequences of approximately 12,000 base pairs (bps), limiting their ability to capture long-range dependencies. Furthermore, the use of k-mer tokenizers reduces modeling resolution, posing challenges for tasks like Single Nucleotide Polymorphism (SNP) analysis.
State Space Models
State Space Models (SSMs) have been applied to genomic tasks, with HyenaDNA [15] utilizing the Hyena operator [37] to process sequences up to 1 million base pairs (bps). Despite this capability, its unidirectional design limits the model’s ability to capture bidirectional genomic contexts [26]. To overcome this, Caduceus [8] introduces a bidirectional Mamba architecture, which aggregates information from both upstream and downstream sequences, enhancing genomic context comprehension. However, SSM-based models often face challenges in memory recall tasks when compared to transformer-based approaches [38]. Additionally, their masking-based training paradigm is less efficient, as it uses only 15% of the data for loss computation per iteration. In contrast, autoregressive training paradigms take advantage of nearly 99% of the data, significantly improving training efficiency and overall performance. Latest SSM-based DNA models, such as Evo [16] and Evo2 [17], introduce stripedHyena and StripedHyena2, showing better scaling rate and improved throughput compared to HyenaDNA. However, to gain better generative performance, they still rely on the autoregressive training paradigm, facing the same limitation as HyenaDNA for bidirectional understanding.
Hybrid Models
Hybrid models incorporate multiple encoding mechanisms such as convolutional neural networks (CNNs), Mamba, and Attention to leverage the complementary strengths of each architecture. Enformer [36] combines CNNs with Transformers, enabling the model to capture long-range genomic interactions spanning up to 100 kilobases. HybriDNA [39] integrates Mamba and Transformer components, extending its receptive field to 131 kilobases.
Appendix B Experimental Details
Appendix C Ablation Experiments
To determine the optimal hybrid architecture for DNA sequence modeling that effectively balances local and global attention, we perform ablation experiments on various configurations of the unidirectional encoder (i.e., modules preceding the bidirectional fusion layer). Referring to [21], we also explore the value of the additional mid-attention layer. Specifically, we evaluate the following configurations: 1) mamba and FFN blocks only, 2) mamba and FFN blocks with mid-attention, 3) mamba and FFN blocks with MoE, and 4) mamba and FFN blocks with both mid-attention and MoE. The ratio of MoE to replace FFNs is set to 0.5. In models with mid-attention, the two independent bidirectional Mamba blocks at the 4th layer are replaced with two independent causal Attention blocks implemented using FlexAttention [23].
The models are pre-trained on sequences of lengths 1024 and 131072, with batch sizes of 128 and 1, respectively, using a single GPU. All other hyperparameters and training settings are consistent with the pre-training setup described earlier. The training perplexity results for these configurations are shown in Figure 4.
Figure 4: Training perplexity of Mid-attention and MoE ablation models on 1024-length and 131k-length sequences.
The training perplexity results reveal that the model with mamba, FFN, and mid-attention exhibits higher perplexity compared to the model with only mamba and FFN, while the model with mamba, FFN, and MoE achieves lower perplexity. This suggests that mid-attention may not enhance training efficiency, whereas MoE contributes positively. Notably, the model combining mamba, FFN, mid-attention, and MoE achieves the lowest perplexity, indicating that mid-attention and MoE can synergistically improve training performance.
To further investigate the role of mid-attention, we conduct additional ablation experiments comparing models with and without mid-attention (all incorporating MoE due to its demonstrated benefits) on the Nucleotide Transformer Benchmark with 3-fold cross-validation since Genomic Benchmark is too saturated to show apparent differences. We pretrain and fine-tune models with hidden dimensions of 32, 72, and 128. The results, summarized in Figure
5, indicate that models with mid-attention perform better at a hidden dimension of 32. However, as the hidden dimension increases, models without mid-attention outperform those with mid-attention at a dimension of 72. At a dimension of 128, the performance of models with and without mid-attention becomes comparable. These findings suggest that mid-attention provides diminishing benefits as the hidden dimension grows larger, eventually becoming negligible at higher dimensions.
Figure 5: Mid attention ablation results on Nucleotide Transformer.
Given that larger hidden dimensions are generally preferred for large-scale DNA sequence modeling, and considering that mid-attention introduces additional computational overhead, we prefer to use models with mamba and FFN blocks without mid-attention for downstream tasks. Nonetheless, we include the experiments of the model with mid-attention in our formal evaluations on Genomic Benchmark and Nucleotide Transformer Benchmark to ensure completeness and provide a comprehensive analysis.
Appendix D Formal Experiment
D.1 Pre-training
Architecture configuration
We utilize Mamba, FFN, MoE blocks as the primary building blocks for unidirectional representation, which are then followed by a bidirectional fusion layer. There are 8 layers in each unidirectional encoder and each layer consists of one Mamba block and one FFN block. The MoE block is to replace the FFN block at a certain ratio, which is set to 0.5 in our experiments. The number of experts is set to 16 and the dimension of FFN is set to 4 times the hidden dimension. The bidirectional fusion layer is achieved by a FlexAttention layer [23] with 4 attention heads.
Meanwhile, we also implement a version of the model with mid-attention, which replaces the Mamba block at the fifth layer with a mid-attention layer. The mid-attention layer is implemented with FlexAttention with 4 attention heads. The attention mask is set to half triangle to allow the model to only attend to the tokens ahead of the current token to keep causality.
Pre-training setup
To ensure a fair comparison with prior work, we pre-train our model on the human reference genome (HG38 [29]) following the training setup described in [8]. We use cross-entropy loss for pre-training. The model is trained with a learning rate of $8\times 10^{-3}$, maintaining a constant token count of $2^{20}$ tokens per batch. Two sequence lengths are used: 1024 and 131072, with corresponding batch sizes of 128 and 1, respectively, across 8 GPUs. Optimization is performed using AdamW [40] with a weight decay of 0.1, $\beta_{1}=0.9$, and $\beta_{2}=0.95$. A cosine learning rate scheduler is applied, incorporating a warmup phase for 10% of the training steps. The learning rate starts at $1\times 10^{-6}$ and peaks at $1\times 10^{-4}$. The coefficient for the MoE auxiliary loss is set to 0.2. The gradient clipping threshold is set to 1.0.
For the 1024-length model, training is conducted for 10,000 steps, while the 131072-length model is trained for 50,000 steps.
We pre-trained three scales of the model: 32, 72, and 144 hidden dimensions to keep the same activated parameters for fair comparison with baseline models on different benchmarks. The hidden dimensions of 32 and 72 are used for the 1024-length model, while the 144 hidden dimension is used for the 131072-length model. The feedforward dimension is set to 4 times the hidden dimension, and the number of attention heads is set to 8. The coefficient for the MoE auxiliary loss is set to 0.2, the number of experts is set to 16. All hyperparameter settings are listed as Table 4.
Table 4: Hyperparameter settings for $\sf JanusDNA$ pretraining (select models).
JanusDNA
w/ Midattn w/o midattn
Layers 8 8 8 8 8
Width 32 72 32 72 144
activated Params (M) 0.422 1.980 0.428 1.989 7.664
total Params (M) 1.798 8.947 1.804 8.956 35.533
Max seq. len. 1024 1024 1024 1024 131072
Batch size 1024 1024 1024 1024 8
global steps 10k 10k 10k 10k 50k
Expert number of MoE 16 16 16 16 16
head number of Attention 4 4 4 4 4
multiple number of FFN width 4 4 4 4 4
coefficient of Auxiliary MoE loss 0.2 0.2 0.2 0.2 0.2
runtime(H800 with evaluation every 2k steps) 3h3m 3h7m 3h2m 3h8m 9h17m
Optimizer AdamW
Optimizer momentum $\beta_{1}$, $\beta_{2}$ = 0.9, 0.95
Learning rate $8\mathrm{e}^{-3}$
LR Scheduler Cosine decay
Weight decay (model) 0.1
D.2 Benchmarks
Genomic Benchmark
For the Genomic Benchmark tasks, we follow the experimental setup of [8] and report their results for comparison. To ensure a fair evaluation, we fine-tune 32-dimensional models to match the activated parameter count of the baselines.
We apply 5-fold cross-validation, splitting the training set into 90/10 train/validation splits and using early stopping based on validation performance with seeds of $\{1,2,3,4,5\}$. Models are fine-tuned for 10 epochs with a batch size of 256.
For learning rate selection, we perform hyperparameter tuning over ${1\times 10^{-3},2\times 10^{-3}}$ as [8], and report the best-performing configuration across cross-validation, as summarized in Table 5. We use cross-entropy loss for fine-tuning. For JanusDNA, the coefficient for the MoE auxiliary loss is set to 0.2.
Table 5: $\sf JanusDNA$ models with and without mid-attention hyperparameter selection for learning rate on genomic benchmarks for top-1 accuracy averaged over 5-fold cross-validation.
Dataset w/ midattn w/o midattn
Mouse Enhancers $1\mathrm{e}^{-3}$ $2\mathrm{e}^{-3}$
Coding vs. Intergenomic $1\mathrm{e}^{-3}$ $1\mathrm{e}^{-3}$
Human vs. Worm $1\mathrm{e}^{-3}$ $1\mathrm{e}^{-3}$
Human Enhancers Cohn $1\mathrm{e}^{-3}$ $1\mathrm{e}^{-3}$
Human Enhancer Ensembl $1\mathrm{e}^{-3}$ $1\mathrm{e}^{-3}$
Human Regulatory $1\mathrm{e}^{-3}$ $1\mathrm{e}^{-3}$
Human OCR Ensembl $1\mathrm{e}^{-3}$ $2\mathrm{e}^{-3}$
Human NonTATA Promoters $1\mathrm{e}^{-3}$ $1\mathrm{e}^{-3}$
Nucleotide Transformer Tasks
For the Nucleotide Transformer tasks, we adopt the experimental setup from [8] and report their results for comparison. To ensure a fair comparison, we fine-tune 72-dimensional models to match the activated parameter count of the baseline models.
We use 10-fold cross-validation with seeds of $\{1,2,3,4,5,6,7,8,9,10\}$, splitting the dataset into 90/10 train/validation subsets and applying early stopping based on validation performance. Each model is fine-tuned for 20 epochs.
We use cross-entropy loss for fine-tuning. We conduct hyperparameter tuning over the following search space, the same as [8]: learning rates in ${1\times 10^{-3},2\times 10^{-3}}$ and batch sizes in ${128,256,512}$. For each task, we report the best-performing configuration across cross-validation, as summarized in Table 6. For JanusDNA, the coefficient for the MoE auxiliary loss is set to 0.2.
Table 6: $\sf JanusDNA$ Hyperparameter Selection for Nucleotide Transformer Tasks. Model with Mid-Attention and Model without Mid-Attention fine-tuning hyperparameters chosen based on best performance averaged over 10-fold cross-validation.
JanusDNA
w/ Midattn w/o midattn
LR batch size LR batch size
Histone markers H3 $1\mathrm{e}^{-3}$ 256 $2\mathrm{e}^{-3}$ 128
H3k14ac $1\mathrm{e}^{-3}$ 256 $1\mathrm{e}^{-3}$ 256
H3k36me3 $1\mathrm{e}^{-3}$ 256 $2\mathrm{e}^{-3}$ 512
H3k4me1 $1\mathrm{e}^{-3}$ 256 $1\mathrm{e}^{-3}$ 256
H3k4me2 $1\mathrm{e}^{-3}$ 256 $1\mathrm{e}^{-3}$ 256
H3k4me3 $1\mathrm{e}^{-3}$ 256 $1\mathrm{e}^{-3}$ 256
H3k79me3 $1\mathrm{e}^{-3}$ 256 $1\mathrm{e}^{-3}$ 256
H3K9ac $1\mathrm{e}^{-3}$ 256 $1\mathrm{e}^{-3}$ 128
H4 $1\mathrm{e}^{-3}$ 256 $1\mathrm{e}^{-3}$ 256
H4ac $1\mathrm{e}^{-3}$ 256 $1\mathrm{e}^{-3}$ 256
Regulatory annotation Enhancers $1\mathrm{e}^{-3}$ 512 $1\mathrm{e}^{-3}$ 512
Enhancers types $2\mathrm{e}^{-3}$ 256 $2\mathrm{e}^{-3}$ 256
Promoter all $1\mathrm{e}^{-3}$ 512 $1\mathrm{e}^{-3}$ 128
Promoter no tata $1\mathrm{e}^{-3}$ 256 $1\mathrm{e}^{-3}$ 128
Promoter tata $2\mathrm{e}^{-3}$ 256 $1\mathrm{e}^{-3}$ 128
Splice site annotation Splice sites acceptors $2\mathrm{e}^{-3}$ 128 $2\mathrm{e}^{-3}$ 128
Splice sites all $2\mathrm{e}^{-3}$ 128 $2\mathrm{e}^{-3}$ 128
Splice sites donors $2\mathrm{e}^{-3}$ 128 $2\mathrm{e}^{-3}$ 128
DNALONGBENCH
For the DNALONGBENCH eQTL tasks, we compare JanusDNA with both the expert model Enformer and the state-of-the-art architecture Caduceus-PH. We obtain the pre-trained weights of Caduceus-PH from Hugging Face: https://huggingface.co/kuleshov-group/caduceus-ph_seqlen-1k_d_model-256_n_layer-4_lr-8e-3.
For all models, we extract the hidden state embeddings from the final layer and apply a pooling layer to obtain a fixed-length representation for each input sequence. A linear classification head is then used to map these representations to the target number of classes for each cell type.
To ensure comparability, we fine-tune the JanusDNA model with 144-dimensional embeddings, matching the activated parameter count of Caduceus-PH. Fine-tuning is conducted for 3 epochs using a learning rate of $4\times 10^{-4}$ and a batch size of 8. We use cross-entropy loss for fine-tuning, and for JanusDNA, the coefficient for the MoE auxiliary loss is set to 0.02. Training is distributed across eight 80GB GPUs, with each GPU processing one batch. All models are fine-tuned and evaluated using float32 precision to ensure stability and fairness in comparison. Due to limited computational resources, we conduct a single run per sub-dataset using the same set of hyperparameters across all experiments.
Appendix E Details of resources used
We use 80GB NVIDIA H100, A100, A800 GPUs for pre-training and fine-tuning.